OpenStudy (anonymous):

A point is moving along the circle x2 + (y − 2)2 = 2 in such a way that its x-coordinate is decreasing at a rate of 6 unit per second. What is happening to the y-coordinate at the instant when the point has reached (−1, 3)

OpenStudy (amistre64):

if we take the derivative implicitly for time, and plug in the values we get the results

OpenStudy (amistre64):

2xx' + 2y'(y-2) = 0 xx' + y'(y-2) = 0 y'(y-2) = -xx' y' = -xx'/(y-2) we are told x' = -6, x=-1, y=3